Exercise 14-42 (Algo) Effects of Current Cost on Performance Measurements (LO 14-2, 5)

Upper Division of Lower Company acquired an asset with a cost of $580,000 and a four-year life. The cash flows from the asset, considering the effects of inflation, were scheduled as follows.

 

YearCash Flow
1$185,000
2
240,000
3
285,000
4
310,000

 

The cost of the asset is expected to increase at a rate of 20 percent per year, compounded each year. Performance measures are based on beginning-of-year gross book values for the investment base. Ignore taxes.

 

Required:

a. What is the ROI for each year of the asset's life, using a historical cost approach?

b. What is the ROI for each year of the asset's life if both the investment base and depreciation are determined by the current cost of the asset at the start of each year?
Answer #1

A) step 1 580,000 /4= 145,000

year 1) 185,000- 145,000= 40,000/ 580,000= 0.0689655172 or Answer: 6.9%

year 2) 240,000- 145,000= 95,000/ 580,000= 0.1637931034 or answer: 16.4%

year 3) 285,000- 145,000= 140,000/ 580,000= 0.2413793103 or answer: 24.1%

year 4) 310,000- 145,000= 165,000/ 580,000= 0.2844827586 or answer: 28.4%


B)

year 1) (185- 145)/ 580= 0.689655172 or answer: 6.9%


1+ 20%= 1.2

year 2) (240- 1/4* 580*1.2)/ 580* 1.2

              240- 145* 1.2/ 696

                (240- 174)/ 696

                     66/ 696= 0.0948275862 or answer: 9.5%


year 3) (285- 1/4 *580* (1.2)^2)/ 580* (1.2)^2

               285- 145* (1.2)^2/ 835.2

                       76.2/ 835.2= 0.0912356322 or Answer 9.1%


year 4) 310- 1/4 *580* (1.2)^3/ 580* (1.2)^3

               310- 145* (1.2)^3/ 1,002.24

                         59.44/ 1,002.24= 0.059307152 or answer: 5.9%